OpenStudy (anonymous):

I've ben asking my question for HOURS!!! Can you PLEASE see if you can solve it and show work. Given the demand function: Qd = 2000 - 3p +.10I +10T (average values of I and T of 1000 and 5 respectively) Please show work!

OpenStudy (dumbcow):

Q = 2150 - 3p solving for p p = (2150-Q)/3 what else is there to solve, whats the rest of the question??